OpenStudy (anonymous):

MEDAL WILL BE GIVEN-divide the two expressions, 3x-6 and x-2

OpenStudy (ivettef365):

you can simplify 3x -6 and you have 3(x-2) ------- x-2

OpenStudy (ivettef365):

now you can cancel x-2

OpenStudy (anonymous):

\[\frac{ 3x-6}{x-2} = 3\frac{(x-2)}{(x-2)} =3\]